OpenStudy (anonymous):

solve 0.002(7-k)+0.04(k-2)=1

OpenStudy (anonymous):

Hint: 0.002 = \(\frac{1}{500}\) 0.04 = \(\frac{1}{25}\)

OpenStudy (anonymous):

Hint 2: \[\frac{19}{500}k-\frac{33}{500}=1\] Any more hints and I'd have to give away the answer ;-)